OpenStudy (anonymous):

FAN AND MEDAL A candy jar contains candies of the same size in three flavors: lime, cherry, and licorice. If the probability of randomly pulling out a lime candy is 3/8, and the probability of randomly pulling out a cherry candy is 2/8, what is the probability of randomly pulling out a licorice flavored candy?

OpenStudy (anonymous):

@danicap @GreatGuy @just_one_last_goodbye

OpenStudy (anonymous):

There are 8 candies total. 3 of them are lime, 2 of them are cherry, and the rest are licorice. So 3+2+l=8. Can you solve for l?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

3+2+5=8? No, l will not be 5. Just get l alone. l=8-3-2

OpenStudy (anonymous):

wait sorry i read it wrong it is 3

OpenStudy (anonymous):

Right. So 3 licorice out of 8 total. Now just write 3 out of 8 in fraction form and you're done.

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

3+2=5 so its 5/8 so licorice would be 3/8

OpenStudy (anonymous):

No problem :)